# DiP: A Scalable, Energy-Efficient Systolic Array for Matrix Multiplication Acceleration

### v1

@@ Abstract
Transformers are gaining increasing attention across Natural Language Processing (NLP) application domains due to their outstanding accuracy. However, these data-intensive models add significant performance demands to the existing computing architectures. Systolic arrays are spatial architectures that have been adopted by commercial AI computing platforms (like Google TPUs), due to their energy-efficient approach of data-reusability. However, these spatial architectures face a penalty in throughput and energy efficiency due to the need for input and output synchronization using First-In-First-Out (FIFO) buffers.
This paper proposes a novel scalable systolic-array architecture featuring Ḏiagonal-I̱nput and P̱ermutated weight-stationary (DiP) dataflow for the acceleration of matrix multiplication. The proposed architecture eliminates the synchronization FIFOs required by state-of-the-art weight stationary systolic arrays. Aside from the area, power, and energy savings achieved by eliminating these FIFOs, DiP architecture maximizes the computational resources (PEs) utilization. Thus, it outperforms the weight-stationary counterparts in terms of throughput by up to 50%. Analytical models are developed for both weight stationary and DiP architectures, including latency, throughput, time to full PEs utilization, and FIFOs overhead. Additionally, a comprehensive hardware design space exploration is demonstrated using commercial 22nm technology, highlighting the scalability advantages of DiP over the conventional approach across various dimensions where DiP offers improvement of energy efficiency per area up to 2.02x. Furthermore, DiP is evaluated using various transformer workloads from widely-used models, consistently outperforming TPU-like architectures, achieving energy improvements of up to 1.81x and latency improvements of up to 1.49x across a range of transformer workloads. At a 64x64 size with 4096 PEs, DiP achieves a peak performance of 8.2 TOPS with energy efficiency 9.55 TOPS/W.
@@ I Introduction
Artificial Intelligence (AI) is continuously dominating various application domains that are vital in our daily life[1]. Natural Language Processing (NLP) is one of the emerging AI applications that are gaining an increasing attention nowadays [2, 3]. Thanks to Transformers, NLP tasks have been revolutionized by providing highly effective and scalable models for language understanding and generation [4]. The exceptional capabilities of these new NLP models have led to a transformer-driven transformation across numerous application domains, including Machine Translation [5], Speech Recognition [6], Multimodal Applications [7], and Computer Vision [8].
@@
However, transformers are data-intensive models that handle massive workloads in comparison to Deep Neural Networks (DNNs) and Convolutional Neural Networks (CNNs) [9]. Additionally, transformer models have been growing exponentially, evolving from the original vanilla transformer model with around 65 million parameters to models with hundreds of billions of parameters [10, 11]. A clear example of the new challenging level of scalability is GPT (Generative Pre-trained Transformer) model (the core of ChatGPT) that incorporates billions of parameters, primarily involving matrix multiplications [12].
@@
Conventional Von-Neumann architectures are struggling to meet these increasing performance demands due to the memory/data-movement bottleneck. Systolic arrays, introduced in 1970s, are spatial architectures that aim at maximizing the data utilization to mitigate the memory/data-movement bottleneck; These architectures are receiving increased attention nowadays as a promising architecture for AI hardware acceleration [13]. Usually, the systolic array consists of a set of two-dimensional (2D) interconnected Processing Elements (PEs). The PE is composed of basic arithmetic, mainly multiplication and accumulation, along with register units. Systolic arrays are spatial architectures that enhance local data utilization by increasing the number of computation operations per each memory access. Accordingly, data flows among different PEs in a wave fashion, while the communication with the synchronization First-In-First-Out (FIFO)s occurs only at the boundary PEs. Moreover, the interconnection of the systolic array naturally realizes data reuse through the interchange of data via PEs, especially in matrix multiplications [14].
@@
Although many systolic arrays are extensively used for hardware accelerators [15, 16, 17, 18, 19, 20, 21], different adopted dataflows require further synchronization interfacing hardware that limits not only the energy efficiency, but also the performance capabilities of these dataflows. Therefore, transformers with massive matrix multiplication workloads will have serious challenges to leverage the systolic array scalability for a next generation of AI hardware.
@@
Tensor Processing Unit (TPU) is one of the well-known AI computing architectures that is introduced by Google to handle massive matrix multiplication workloads with higher performance and energy efficiency than CPUs and GPUs [22]. TPUs adopt weight stationary dataflow, which maximizes the data utilization of both weights and inputs. The first-generation TPU (TPU v1) is designed primarily for inference, featuring a 256x256 systolic array optimized for 8-bit integer (INT8) operations, achieving a peak throughput of 92 TOPS [23]. TPU v2 shifted to mixed-precision training with a smaller 128x128 systolic array per core optimized for FP16 and bfloat16 operations, boosting throughput to 180 TeraFLOPS [24]. TPU v3 and v4 maintained the same array size but v3 doubled the throughput to 420 TFLOPS per chip, aided by high memory bandwidth, while TPU v4 uses four cores of 128x128 architecture, achieving up to 1 PFLOPS [25]. In conventional WS systolic array, synchronization FIFOs are necessary to synchronize both inputs and outputs, adding significant overhead in terms of throughput, energy, latency, power, and area. Moreover, the computations propagate as a diagonal wavefront from the top-left corner to the bottom-right corner of the systolic array due to the WS dataflow. This significantly decreases the overall PEs utilization resulting in degraded performance and increased latency.
@@
Meissa is one of systolic architectures that separates multipliers from the adders rather than combining them in a unified array [26]. It stands for multiplying matrices efficiently In a scalable systolic architecture. like to the TPUs, it adopted the WS dataflow for the systolic array, but it eliminates the input synchronization FIFOs to reduce the overall latency. However, it has bulky adder trees per each column of the systolic array instead of having the partial summations accumulated through the PEs. These adder trees impose scalability limitations due to serious physical implementation challenges. The larger the adder trees the deeper pipelines they require to achieve higher frequency. This increases the overall latency, area, and energy consumption. The routing congestion is another expensive challenge, caused by delivering all products from all PEs in the same column to the adder tree. Consequently, Meissa is not scalable to large NxN dimensions, which is vital to large language models. Moreover, it still requires the output synchronization FIFOs which still add a considerable area/power/energy penalty.
@@
In this paper, we present a novel Diagonal-Input Permutated (DiP) weight-stationary Systolic Array that overcomes the main challenges of the conventional WS systolic arrays. DiP is a scalable architecture of NxN PEs that maximizes the PEs utilization by featuring diagonal-input movement and permutated weight-stationary dataflow for the acceleration of matrix multiplication.
The main contributions of this work are highlighted as follows:
@@
  • We introduce DiP, a novel scalable spatial architecture that maximizes the PEs utilization and energy efficiency achieving improvement in throughput by up to 1.49x and energy efficiency per area by up to 2.017x.
@@
  • The proposed architecture eliminates the input/output synchronization FIFOs of the WS by implementing a new dataflow with diagonal-input movement and permutated weights.
@@
  • The analytical models for DiP and WS including throughput, latency, time to full PEs utilization (TFPU), and registers overhead are extracted for different systolic array sizes.
@@
  • Hardware Design space exploration and implementation are presented for DiP and WS using commercial 22nm technology, offering energy efficiency per area improvement up to 2.02x with area and power consumption savings up to 8.12% and 19.95%, respectively.
@@
  • DiP is evaluated using various transformers workloads from widely-used models, outperforming TPU-like architectures, and achieving energy improvement up to 1.81x and latency improvement up to to 1.49x across various transformer workloads.
@@ II Systolic Arrays Background
Systolic arrays usually adopt one of the dataflows to control the data movement across the PEs. Each dataflow retains one of the input/output/weight data to be stationary during computations for the maximum duration to exploit the data reuse [27, 28]. The following dataflows are the most common for the systolic array design:
  • Weight Stationary (WS): weights are initially loaded to PEs and kept stationary during processing. The input matrix and partial summations are moved among PEs during processing.
  • Input Stationary (IS): input matrix is initially loaded to the systolic array, while weight matrix and partial summations are moved among PEs during processing.
  • Output Stationary (OS): input and weight matrices are moved across the PEs, while the partial summations are accumulated inside the PEs.
  • Row Stationary (RS): This dataflow is proposed by Eyeriss [29]. It adopts spatial architecture that uses coarse-grained PEs with internal memories to store weights and inputs. Inputs are broadcasted diagonally across the PEs, while weight matrix broadcasted horizontally, and partial summations move vertically.
OS dataflow moves both input and weight matrices simultaneously, which effectively doubles the required memory bandwidth for the systolic array. With RS dataflow, data redundancy increases because copies of the data are loaded into different PEs. Additionally, the circulation of weights within each PE reduces energy efficiency. On the other hand, the WS dataflow is widely used in many architectures, such as the Google TPU, due to its scalability and flexibility in handling convolutions and matrix multiplication [23, 24, 25]. Additionally, it requires less memory bandwidth. Therefore, we will focus on improving it.
@@ II-A WS Dataflow
WS dataflow is widely used in many architectures where weights are initially loaded to the systolic array’s PEs, while input matrix circulates among the PEs in a systolic fashion. This approach mitigates the memory bottleneck by reducing the number of memory accesses and increasing data reusability. However, the WS dataflow requires input and output FIFOs to synchronize the data for proper functionality. Figure 1 shows NxN weight stationary systolic array. There are two FIFO groups for input/output synchronization. The input FIFO group consists of series of FIFOs with incrementally increasing depth starting from one element in the second row up to $N-1$ elements in the last row. The output FIFO group is structured similarly to the input FIFO group, but the FIFO depths decrease from $N-1$ elements to one element moving from left to right across the systolic array columns. Weights are loaded vertically, and partial summations (psums) are accumulated vertically. Input matrix is shifted horizontally from the input FIFO group, and the output from the last PE row is shifted out to the output FIFO group. The use of FIFOs increases area, power, energy consumption, and latency for matrix multiplication. Consequently, FIFO-based systolic arrays suffer from lower energy efficiency, and reduced PE utilization, leading to higher latency and decreased throughput.
Figure 1: Top-level schematic for NxN weight stationary systolic array. There are two FIFO groups for input/output synchronization. The weights (black crossed buses) are loaded vertically, and psums (grey buses) are accumulated vertically. Inputs (I) (blue buses) are shifted horizontally from the input FIFO group, and the output is shifted out to the output FIFO group.
@@
The analytical modeling of the WS systolic array is studied for latency, throughput, TFPU, and FIFOs overhead, providing insights into dataflow performance. For the WS latency analytical model, as shown in (1), WS requires $33N+S-$ 3 cycles to complete the processing, where $N$ represents the number of rows/columns per WS systolic array, and $S$ is the number of pipeline stages per Multiply-Accumulate (MAC) unit, equals one for 1-stage pipelined MAC, and two for 2-stage pipelined MAC. The throughput, as indicated in (2), is calculated as the ratio of total operations to WS latency. Regarding register overhead, WS systolic array uses two FIFO groups for input and output synchronization. Each group consists of $N-1$ FIFOs, as shown in Fig. 1. Additionally, each FIFO group includes ${N(N-1)/2}$ registers. Consequently, the total register overhead for a typical WS systolic array is calculated as shown in (3). TFPU is another metric introduced to calculate the required number of cycles to reach full utilization of PEs. This metric shows the overhead when the input matrix is initially loaded to PEs. TFPU is calculated, as shown in (4), where it takes $12N-$ 1 cycles for WS to reach full PEs utilization.
@@ III DiP Architecture
Figure 2: (a) General NxN DiP systolic array architecture, inputs(I) move diagonally across PE rows, transitioning from one row to the next. The boundary PEs are diagonally connected, so that the registered inputs from the leftmost PE column feed into the inputs of the rightmost PE column in the subsequent row. Weights are loaded vertically, and psums are accumulated vertically along the columns as well. (b) PE block diagram, consisting of 2-stage pipelined MAC unit and four enabled registers. Control signals wshift, pe_en, mul_en, and adder_en are used for operations control. wshift is shared between all systolic array PEs, while pe_en, mul_en, and adder_en are shared across each PE row. The grey buses represent the partial sum (psum) buses, while the weight buses are indicated by black crossed buses. Additionally, the input data buses are shown in blue, and control signals are represented with dashed lines. (c) General weight matrix permutation for DiP dataflow. The weight matrix is permutated by shifting and rotating each column by its column index
Figure 3: Pseudocode for weight matrix permutation, where each column undergoes an incremental row shift based on the column index, creating a unique, wrap-around pattern across rows.
Figure 4: Example for 3x3 DiP systolic array (a) shows the diagonal input connections for 3x3 DiP, (b) shows 3x3 DiP weight matrix permutation by shifting and rotating each column by its column index, and (c) shows the processing flow for 3x3 DiP example cycle by cycle. Cycles (-2, -1, 0) are dedicated to weight matrix loading, while Cycle 0 involves loading the last row of the weight matrix and the first row of the input matrix. Cycles from Cycle 1 to Cycle 5 are allocated for matrix multiplication processing, with final output rows becoming ready starting from Cycle 3.

### v2

@@ Abstract
Transformers are gaining increasing attention across Natural Language Processing (NLP) application domains due to their outstanding accuracy. However, these data-intensive models add significant performance demands to the existing computing architectures. Systolic array architectures, adopted by commercial AI computing platforms like Google TPUs, offer energy-efficient data reuse but face throughput and energy penalties due to input-output synchronization via First-In-First-Out (FIFO) buffers.
This paper proposes a novel scalable systolic array architecture featuring Ḏiagonal-I̱nput and P̱ermutated weight stationary (DiP) dataflow for matrix multiplication acceleration. The proposed architecture eliminates the synchronization FIFOs required by state-of-the-art weight stationary systolic arrays. Beyond the area, power, and energy savings achieved by eliminating these FIFOs, DiP architecture maximizes the computational resource utilization, achieving up to 50% throughput improvement over conventional weight stationary architectures. Analytical models are developed for both weight stationary and DiP architectures, including latency, throughput, time to full PEs utilization (TFPU), and FIFOs overhead. A comprehensive hardware design space exploration using 22nm commercial technology demonstrates DiP’s scalability advantages, achieving up to a 2.02× improvement in energy efficiency per area. Furthermore, DiP outperforms TPU-like architectures on transformer workloads from widely-used models, delivering energy improvement up to 1.81× and latency improvement up to 1.49×. At a 64×64 size with 4096 PEs, DiP achieves a peak throughput of 8.192 TOPS with energy efficiency 9.548 TOPS/W.
@@ I Introduction
Artificial Intelligence (AI) is increasingly dominating various application domains [1]. Natural Language Processing (NLP) is one of the crucial AI application domains receiving increasing attention nowadays [2, 3]. Thanks to Transformers, NLP tasks have been revolutionized by providing highly effective and scalable models for language understanding and generation [4]. The exceptional capabilities of these models have led to a transformer-driven transformation across numerous application domains, including Machine Translation [5], Speech Recognition [6], Multimodal Applications [7], and Computer Vision [8].
@@
However, transformers are data-intensive models that handle massive workloads compared to Deep Neural Networks (DNNs) and Convolutional Neural Networks (CNNs) [9]. Additionally, transformer models have been growing exponentially, evolving from the original vanilla Transformer model with around 65 million parameters to models with hundreds of billions of parameters [10, 11]. A clear example of the challenging level of scalability is GPT (Generative Pre-trained Transformer) model that incorporates billions of parameters, primarily involving matrix multiplications [12].
@@
Conventional Von-Neumann architectures are struggling to meet these increasing performance demands due to the memory/data-movement bottleneck. Systolic arrays, introduced in 1982, are spatial architectures that designed to maximize the data utilization to mitigate the memory bottleneck. These architectures are receiving increased attention nowadays as a promising architecture for AI hardware acceleration [13]. Typically, a systolic array consists of a two-dimensional (2D) interconnected Processing Elements (PEs). The PE consists of basic arithmetic, mainly multiplication and accumulation, along with register units. Systolic arrays are spatial architectures that enhance local data utilization by increasing the number of computation operations per each memory access. Accordingly, input data circulates among PEs in a wave-like fashion, while the communication with the synchronization First-In-First-Out (FIFO)s occurs only at the boundary PEs. Moreover, the interconnection of the systolic array naturally facilitates data reusability among PEs, particularly during matrix multiplications [14].
@@
Although many systolic arrays are used for hardware acceleration [15, 16, 17, 18, 19, 20, 21], different adopted dataflows require additional synchronization hardware, which limits the energy efficiency and constrains performance capabilities. Therefore, transformers with massive matrix multiplication workloads will face significant challenges in leveraging the systolic arrays for the next generation of AI hardware.
@@
Tensor Processing Unit (TPU) is a well-known AI computing architectures introduced by Google to handle massive matrix multiplication workloads with higher performance and energy efficiency than CPUs and GPUs [22]. TPUs adopt weight stationary (WS) dataflow, which maximizes the data utilization of both weights and inputs. Google TPU V1 is designed primarily for inference, featuring a 256×256 systolic array optimized for 8-bit integer (INT8) operations, achieving a peak throughput of 92 TOPS [23]. TPU V2 shifted to mixed-precision training with a smaller 128×128 systolic array per core optimized for FP16 and bfloat16 operations, boosting throughput to 180 TeraFLOPS [24]. TPU V3 and V4 maintained the same array size but TPU V3 doubled the throughput to 420 TFLOPS per chip, aided by high memory bandwidth, while TPU v4 integrates four cores of 128×128 architecture, achieving up to 1 PFLOPS [25].
@@
Meissa is a systolic architecture that separates multipliers from adders instead of integrating them into a unified array [26]. Similar to TPUs, it employs WS dataflow for the systolic array but eliminates input synchronization FIFOs to reduce overall latency. However, it has bulky adder trees per each column of the systolic array. These adder trees impose scalability limitations due to serious physical implementation challenges. As adder trees become larger, they require deeper pipelines to achieve higher frequencies. This increases the overall latency, area, and energy consumption. In addition, routing congestion presents another costly challenge, when all partial products from each PE column are delivered to the adder tree. Consequently, Meissa is not scalable to large N×N dimensions, which is critical for computation-intensive workload. Moreover, it still requires the output synchronization FIFOs, which add additional area, power, and energy penalty.
@@
In this paper, we present a novel Diagonal-Input Permutated (DiP) weight stationary systolic array that overcomes the main challenges of the conventional WS systolic array architectures. The main contributions of this work are highlighted as follows:
@@
  • We introduce DiP, a novel scalable spatial architecture of N×N PEs for matrix multiplication acceleration. DiP maximizes PE utilization and energy efficiency, achieving a throughput improvement of up to 1.49×.
@@
  • The proposed architecture eliminates the input/output synchronization FIFOs in WS by implementing a new dataflow with diagonal-input movement and permutated weights.
@@
  • The analytical models are extracted for DiP and WS including throughput, latency, time to full PE utilization (TFPU), and register overhead for different systolic array sizes.
@@
  • Hardware design space exploration and implementation are presented for DiP and WS using commercial 22nm technology, offering up to 2.02× energy efficiency per area improvement.
@@
  • DiP is evaluated using various transformer workloads from widely used models, outperforming TPU-like architectures and achieving up to 1.81× energy efficiency improvement and up to 1.49× latency improvement.
@@ II Systolic Arrays Background
Systolic arrays adopt one of the dataflows to control the data movement across the PEs. Each dataflow retains one of the input, output, or weight data to be stationary during computations for the maximum duration to maximize the data reuse [27, 28]. The following dataflows are the most common in systolic array design:
  • Weight Stationary (WS): weights are initially loaded to the PEs and kept stationary during processing. The input matrix and partial summations move among the PEs.
  • Input Stationary (IS): input matrix is initially loaded to the systolic array, while weight matrix and partial summations move among the PEs.
  • Output Stationary (OS): input and weight matrices move among the PEs, while the partial summations are accumulated inside PEs.
  • Row Stationary (RS): it is proposed by Eyeriss [29]. It adopts spatial architecture that uses coarse-grained PEs with internal memories to store weights and inputs. Inputs are broadcasted diagonally across the PEs, while weight matrix broadcasted horizontally, and partial summations move vertically.
The OS dataflow moves both input and weight matrices simultaneously, which effectively doubles the required memory bandwidth for the systolic array. In RS dataflow, data redundancy increases because copies of the data are loaded into multiple PEs. Additionally, the circulation of weights within each PE reduces energy efficiency. On the other hand, WS dataflow is widely used in many architectures, such as Google TPU, due to its scalability and flexibility in handling convolutions and matrix multiplication [23, 24, 25]. Moreover, it requires less memory bandwidth. Therefore, we focus on improving this dataflow.
@@ II-A WS Dataflow
The WS dataflow is widely used in many architectures where weights are initially loaded to the PEs, while the input matrix circulates between them in a systolic fashion. This approach mitigates the memory bottleneck by reducing memory accesses and increasing data reuse. However, WS dataflow requires input and output FIFOs to synchronize data for proper functionality. Figure 1 shows the architecture of WS systolic array consisting of N×N PEs and two FIFO groups for input/output synchronization. WS PE consists of 2-stage pipelined MAC unit and four enabled registers for input, weight, multiplier output and adder output. Control signals wshift, pe_en, mul_en, and adder_en manage the PE operations. Specifically, wshift enables the weight register, while pe_en enables the input register. mul_en and adder_en selectively enable their respective registers only during active computation cycles, reducing power consumption during inactive cycles. wshift is shared across each PE row, while pe_en, mul_en, and adder_en are shared across each PE column. The input FIFO group consists of series of FIFOs with incrementally increasing depth starting from one element in the second row to $N-1$ elements in the last row. The output FIFO group is structured similarly to the input FIFO group, but the FIFO depths decrease from $N-1$ elements to one element moving from left to right across the systolic array columns.
The input FIFOs map the input matrix to the PEs in diagonal wave fashion, while the output FIFOs synchronize partial summations along each column [27]. In WS, the input matrix and partial summations (psums) are pipelined to ensure proper processing and improve data reuse. The PE in WS is fine-grained, unlike the coarse-grained PE in Eyeriss, which broadcasts weights and input matrices to the array and stores them in internal memories [29]. Weights are loaded vertically, while the input matrix is shifted horizontally, while the output is shifted to the output FIFO group. Using synchronization FIFOs add significant overheads in terms of area, power, energy consumption, and latency during matrix multiplication. Moreover, the computations propagate as a diagonal wave from the top-left corner to the bottom-right corner of the systolic array. This substantially reduces overall PE utilization resulting in degraded performance and increased latency. As a result, FIFO-based systolic arrays suffer from lower energy efficiency, reduced PE utilization, higher latency, and decreased throughput.
Figure 1: (a) Top-level schematic for the WS systolic array with N×N PEs and two input/output FIFO groups. The weights (black crossed buses) are loaded vertically, and psums (grey buses) are accumulated along the columns. (b) WS PE block diagram, consisting of a 2-stage pipelined MAC unit and four enabled registers. Control signals wshift, pe_en, mul_en, and adder_en are used for operations control. wshift is shared across each PE row, while pe_en, mul_en, and adder_en are shared across each PE column.
@@
The analytical modeling of the WS systolic array is studied for latency, throughput, TFPU, and FIFOs overhead, providing insights into dataflow performance. For the WS latency analytical model, as shown in (1), WS requires $33N+S-$ 3 cycles to complete the processing, where $N$ represents the number of rows/columns per WS systolic array, and $S$ is the number of pipeline stages per Multiply-Accumulate (MAC) unit, which equals one for 1-stage pipelined MAC, and two for 2-stage pipelined MAC. The throughput, as indicated in (2), is calculated as the ratio of total operations to WS latency. Regarding register overhead, WS systolic array uses two FIFO groups for input and output synchronization. Each group consists of $N-1$ FIFOs, as shown in Fig. 1. Additionally, each FIFO group includes ${N(N-1)/2}$ registers. Consequently, the total register overhead for a typical WS systolic array is calculated as shown in (3). TFPU is another metric introduced to calculate the required number of cycles to reach full utilization of PEs. This metric shows the overhead when the input matrix is initially loaded to the PEs. TFPU is calculated, as shown in (4), where it takes $12N-$ 1 cycles for WS to reach full PEs utilization.
@@ III DiP Architecture
Figure 2: (a) N×N DiP systolic array architecture, inputs(I) move diagonally across PE rows, transitioning from one row to the next. The boundary PEs are diagonally connected, so that the registered inputs from the leftmost PE column feed into the inputs of the rightmost PE column in the subsequent row. Weights are loaded vertically, and psums are accumulated vertically along the columns as well. (b) PE block diagram, consisting of a 2-stage pipelined MAC unit and four enabled registers. Control signals wshift, pe_en, mul_en, and adder_en are used for operations control. wshift is shared between all systolic array PEs, while pe_en, mul_en, and adder_en are shared across each PE row. (c) General weight matrix permutation for DiP dataflow. The weight matrix is permutated by shifting and rotating each column by its column index
Figure 3: Example for 3×3 DiP systolic array (a) shows the diagonal input connections for 3×3 DiP, (b) shows 3×3 DiP weight matrix permutation by shifting and rotating each column by its column index, and (c) shows the processing flow for 3×3 DiP example cycle by cycle. Cycles (-2, -1, 0) are dedicated to weight matrix loading, while Cycle 0 involves loading the last row of the weight matrix and the first row of the input matrix. Cycles from Cycle 1 to Cycle 5 are allocated for matrix multiplication processing, while the first output row becomes ready at Cycle 3.